### v1

# Quantum Circuit Reconstruction from Power Side-Channel Attacks on Quantum Computer Controllers

Abstract
The interest in quantum computing has grown rapidly in recent years, and with it grows the importance of securing quantum circuits. A novel type of threat to quantum circuits that dedicated attackers could launch are power trace attacks. To address this threat, this paper presents first formalization and demonstration of using power traces to unlock and steal quantum circuit secrets. With access to power traces, attackers can recover information about the control pulses sent to quantum computers. From the control pulses, the gate level description of the circuits, and eventually the secret algorithms can be reverse engineered. This work demonstrates how and what information could be recovered. This work uses algebraic reconstruction from power traces to realize two new types of single trace attacks: per-channel and total power attacks. The former attack relies on per-channel measurements to perform a brute-force attack to reconstruct the quantum circuits. The latter attack performs a single-trace attack using Mixed-Integer Linear Programming optimization. Through the use of algebraic reconstruction, this work demonstrates that quantum circuit secrets can be stolen with high accuracy. Evaluation on 32 real benchmark quantum circuits shows that our technique is highly effective at reconstructing quantum circuits. The findings not only show the veracity of the potential attacks, but also the need to develop new means to protect quantum circuits from power trace attacks. Throughout this work real control pulse information from real quantum computers is used to demonstrate potential attacks based on simulation of collection of power traces.
keywords:
Quantum Circuits Quantum Computers Side Channel Attacks Power Trace Attack Automated Reasoning Mixed Integer Linear Programming
1 Introduction
Figure 1: Typical operation of a cloud-based quantum computer. Red arrows highlight potential power trace threats.
The interest in quantum computing is growing rapidly and already a large numbers of quantum computers are easily accessible over the internet to researchers and everyday users. Due to the expensive nature of the quantum computing equipment, these computers are currently available as cloud-based systems. For example, IBM Quantum [IBM23], Amazon Braket [Ama23], and Microsoft Azure [Mic23] already provide access to various types of Noisy Intermediate-Scale Quantum (NISQ) devices from different vendors. Remote access makes it easy for different users and companies to run algorithms on real quantum computers without the need to purchase or maintain them. On the other hand, the users have no control over the physical space where the quantum computers are. While the cloud providers may not be themselves malicious, the threat of malicious insiders within data centers or cloud computing facilities is well-known in classical security [SJPBL14]. These malicious insiders may have physical access to the equipment of quantum computers. With access to the quantum computers and the microwave controllers, malicious insiders could leverage physically collected information to steal or leak the quantum circuit secrets.
For classical computers, side-channel attacks of different types are a well-known threat [Sze18]. Two widely studied and analyzed side-channels are timing- and power-based channels. There are also thermal, EM, acoustic, and a variety of other categories of side-channels. In timing side-channel attacks the attacker is trying to learn some secret properties about the circuit by measuring the execution times. Timing attacks are powerful enough in classical computers to break the implementation of standard cryptographic primitives, such as DSA, RSA or Diffie-Hellman [Koc96]. Timing side-channels are easier to exploit as they only require doing timing measurement of the victim. Power side-channel attacks are more convoluted attacks, where the attacker tries to establish a correlation between the power consumption and the operations and data that the circuit executes. Kocher et al. [KJJ99] showed how to reconstruct the encryption keys in the Data Encryption Standard (DES) using a power consumption analysis. Power attacks require physical access to monitor the execution of the target computer.
Power side channels are well studied for classical computers [EPMS23, PR13, WSRW21, ${}^{+}$21, ${}^{+}$20, ABP19, WSW19, BCHC18, BYT17, ${}^{+}$13, EW14, MOPT12]. There are even platforms [OC14] for analyzing power side channels. However, understanding power side channels for quantum computers has not been explored yet, which this work aims to address. We show how a malicious attacker can reconstruct a secret quantum circuit that is being executed, by simply measuring the power consumption.
There is very limited research on understanding power side-channel attacks on quantum computers [XES23b]. One insight about quantum computers is that if the attacker is to perform physical measurements on the qubits during the computation, these measurements would interact with the qubits and destroy their state. However, we observe that each quantum computer is controlled by external hardware such as microwave electronics and controllers. Quantum computers, such as superconducting qubit machines from IBM, Rigetti, or others, use microwave pulses to execute gate operations on qubits. The control pulses are fully classical and could be spied on – which is the target of this work.
This paper shows how by measuring the power consumption of the controller devices sending microwave pulses to a quantum computer, we can recover a potentially secret quantum circuit that the quantum computer executed. We show that anybody with access to power traces of the control pulse generation devices can capture and recover the control information. While this work explores power-based side-channels, the same or similar ideas could apply to electromagnetic (EM) or other types of physical side-channels. This is left as orthogonal work.
1.1 Power Side-Channel Threats to Quantum Computers
Figure 1 shows the operation of today’s cloud-based quantum computers. Remote users submit jobs to the cloud provider, where the job management or similar server dispatches the jobs to particular quantum computers, also called backends on IBM Quantum. Typically the digital instructions are sent to controller logic, such as microwave electronics, which generate the actual analog control signals sent to the quantum computer.
We assume that the classical computer components, e.g., the job management server, are protected from side-channels. Meanwhile, controller electronics of quantum computers, such as arbitrary waveform generators (AWGs) have not been analyzed for potential side-channels before this work. Consequently, we focus on and demonstrate potential new, power side-channel attacks that could be used to extract information about users’ quantum circuits (quantum gates and qubits). Rather than targeting the superconducting qubits themselves (which are isolated in a cryogenic refrigerator), we focus on the controller electronics shown in the middle of Figure 1.
The vulnerabilities in quantum computer controllers encompass more than just gate recovery. An attacker might discern the number of qubits used in a quantum program, a significant concern for algorithms like Quantum Approximate Optimization Algorithm (QAOA) [CK19]. Variational quantum algorithms [${}^{+}$21] are notably sensitive to qubit count and circuit depth. Extracting such hyperparameters can reveal crucial information about the algorithm. Moreover, given the current state of quantum computing where inputs are hard-coded, an attacker could potentially extract sensitive input data from the circuit.
Our primary focus is to protect the intellectual property embedded in quantum programs. Quantum circuits encapsulate both the algorithm and its inputs. An attacker capable of recovering parts of the quantum circuit can misappropriate this intellectual property or sensitive data. This concern amplifies when quantum programs run on external quantum computers. A recent workshop by The National Quantum Coordination Office underscored this issue, emphasizing the role of formal methods in enhancing quantum computing security [Off22]. Our research aligns with this perspective, pioneering the use of formal methods for quantum circuit recovery.
1.2 Lessons from Historical Technological Threats
The significance of vulnerabilities in quantum computer controllers cannot be overstated, considering the ongoing evolution of technological advancements and their inherent risks. For perspective, speculative execution attacks in classical computers were not recognized as threats until 2018 [${}^{+}$20, ${}^{+}$20], despite having been operational since the commercialization of the technology by IBM and Intel in the 1990s. This delay between innovation and the identification of vulnerabilities underscores the necessity for proactive security measures.
In the context of quantum computers, addressing vulnerabilities in their infancy is crucial, even as the field continues to develop. Our research delves into these potential risks, emphasizing the importance of safeguarding quantum computing systems.
Drawing from past lessons, unchecked technological vulnerabilities can culminate in substantial security breaches. As cache attacks emerged long after caches were invented, a similar oversight with quantum computers could be costly. It’s imperative to address these challenges proactively, ensuring quantum computing’s advancement aligns with rigorous security protocols.
1.3 Contributions
Compared to the work by Xu et al. [XES23b], our work provides a formalization of the power side channel attack. Additionally, we also present a novel algebraic reconstruction method for recovery of quantum circuits. There are two reconstruction methods that we introduce; those methods depend on the attackers’ abilities. In the per-channel method we assume that the attacker collects power traces from individual qubit channels. Meanwhile, in our total power single trace method we assume the attacker can only measure the total power trace of all channels.
We have empirically evaluated our approach for 32 benchmark quantum circuits. The evaluation shows that our technique is highly effective at reconstructing quantum circuits. In summary, the paper contributes the following:
  • The first formalization of power side channel attacks on complete reconstruction of quantum circuits from power traces, which is given in Section 4.
  • Demonstration of circuit reconstruction using our new per-channel single trace attack: this attack relies on single-shot per-channel measurements to perform a brute-force attack to reconstruct the quantum circuits, which is given in Section 5.1.
  • Demonstration of circuit reconstruction using our new single-shot total power side-channel attack: this attack relies on a single power side-channel measurement and performs attack using Linear Mixed Integer Real Arithmetic (LIRA) solving and Mixed-Integer Linear Programming (MILP) optimization, which is given in Section 5.2.
  • Details of the evaluation of the attacks on 32 real quantum circuits in the QASMBench [^1] https://github.com/pnnl/QASMBench/ benchmark suite [LSKA22], using control pulse information from real IBM quantum computers, which is given in Section 6.
2 Background
This section provides background on quantum computers and typical quantum computer workflow.
Figure 2: Schematic of a typical qubit drive setup. The local oscillator (LO) generates a low phase-noise microwave carrier signal, and then the wave is modulated in the IQ mixer by I and Q components generated by the arbitrary wave generator (AWG). The pulse is then sent to drive the qubits in the quantum computer. The red line shows the process to collect power traces, which can be exploited by attackers to retrieve information.
2.1 Qubits and Quantum States
The most essential component of quantum computing is the quantum bit, or qubit for short. It is theoretically comparable to the bit used in current classical computing. Like a bit, a qubit has two basis states, which are represented by the bra-ket notation as |$\ket{0}$ and |$\ket{1}$. Whereas a classical bit can only be either 0 or 1, a qubit can be any linear combination of |$\ket{0}$ and |$\ket{1}$ with norm 1. With this notation, a qubit |$\ket{\psi}$ is typically represented as: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, where $\alpha$ and $\beta$ are complex numbers satisfying $|\alpha|^{2}+|\beta|^{2}=1$. Qubits are frequently represented using vectors. For example, |$\ket{0}=[1,0]^{T}$ and |$\ket{1}=[0,1]^{T}$ are two-dimensional vectors that can be used to represent the basis states of a single qubit. The above state |$\ket{\psi}$ can therefore be expressed as |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}=[\alpha,\beta]^{T}$. There are representations that are equivalent for multi-qubit states. For instance, the two-qubit states’ space is made up of the four basis states |$\ket{00}$, |$\ket{01}$, |$\ket{10}$, and |$\ket{11}$. In the space of $n$-qubit states, there are $2^{n}$ basis states that range from |$\ket{0\dots 0}$ to |$\ket{1\dots 1}$, and a $n$-qubit state |$\ket{\phi}$ can be expressed as follows:
|$\ket{\phi}=\sum_{i=0}^{2^{n}-1}a_{i}\ket{i}\text{ where }\sum_{i=0}^{2^{n}-1}|a_{i}|^{2}=1.$
2.2 Quantum Gates
The fundamental quantum operations at the logic-level are quantum gates, which are comparable to classical computing. Quantum algorithms are made up of a series of quantum gates that can convert input qubits into different quantum states. Quantum gates are unitary operations that modify the input qubits, i.e., for a quantum gate $U$ that is applied to a quantum state |$\ket{\psi}$, the quantum state is evolved to |$\ket{\psi}\rightarrow U\ket{\psi}$, and $UU^{\dagger}=U^{\dagger}U=I$. With the vector-matrix representation, $2^{n}\times 2^{n}$ matrices can be utilized to express $n$-qubit quantum gates.
One classical example is the gate that is analogous to the NOT gate in classical computing, the Pauli-X gate, that exchanges the components of |$\ket{0}$ and |$\ket{1}$. Another significant example is the two-qubit CNOT gate, also known as the CX gate, which, if the control qubit is in the state |$\ket{1}$, applies a Pauli-$X$ gate to the target qubit; otherwise, nothing happens. There are some more matrices of quantum gates along with their matrix representations. One thing to keep in mind is that our qubit order is consistent with that in Qiskit [Qis23], where the leftmost qubit is the most important and the rightmost qubit is the least important. As a result, if a different qubit order is used in other studies, the CX gate may have a different matrix representation. Below we show several matrices of quantum gates:
${\tt I}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\ {\tt X}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\ {\tt CX}=\begin{bmatrix}1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\end{bmatrix},\ {\tt RZ(\theta)}=\begin{bmatrix}e^{-i\frac{\theta}{2}}&0\\ 0&e^{i\frac{\theta}{2}}\end{bmatrix},\ {\tt SX}=\frac{1}{2}\begin{bmatrix}1+i&1-i\\ 1-i&1+i\end{bmatrix}$
A small number of quantum gates can be used to approximate any unitary quantum gate within a small error, as shown in the study [DBE95]. One of the crucial configurations of quantum computers is the basis gates, also known as native gates. Different manufacturers or even various versions of quantum computers from same manufacturer may have different native gates. Choice of supporting different types of basis gates is a trade-off between numerous attributes like error rate and efficiency. Our experiments in this study were done on IBM Quantum, and typically, the basis gates are provided by IBM quantum computers are: I, RZ, SX, X, and CX. Prior to being executed on physical quantum computing hardware, quantum gates such as the commonly utilized Hadamard gate need to be broken down into these basis gates.
2.3 Control Pulses
(a) CX pulse
(b) X pulse
(c) SX pulse
Figure 3: Example SX, X, and CX gate control pulses. Figure from [XES23b].
Microwave pulses are typically used to control superconducting qubits. The right control pulses corresponding to each basis gate must be generated and supplied to the quantum computer in order for it to execute each basis gate. Figure 3 displays examples of control pulses for the SX, X, and CX gates. The I gate on IBM Quantum has no effect and is effectively just a delay between pulses. In addition, the RZ gate is a virtual gate without an actual pulse.
Typically, the envelope, frequency, and phase together characterize a pulse. In the case of the superconducting qubit control, the frequency and phase specify the carrier signal that is to be modulated by the lower-frequency envelope signal. The local oscillator (LO) generates the low phase-noise microwave carrier signal. The envelope specifies the shape of the signal that is created by the arbitrary waveform generator (AWG). The envelop signal is mixed with the carrier signal and that is transmitted to the qubit or couplings to drive operation of the quantum computer. Figure 2 displays the standard devices for driving the qubits.
Despite the fact that envelopes can have any design, they are often parameterized by a few preset forms, requiring a minimal number of parameters to specify the envelope. These factors often include duration, which indicates how long the pulse is, amplitude, which indicates how strong the pulse is, and other parameters, which determine the pulse’s structure. For instance, the Derivative Removal by Adiabatic Gate (DRAG) pulse, which is defined by sigma, which specifies how wide or narrow the Gaussian peak is, and beta, which specifies the correction amplitude, as well as the duration and amplitude, is a standard Gaussian pulse with an additional Gaussian derivative component and lifting applied. Another illustration is the Gaussian square pulse, which is a square pulse with a rise-fall in the shape of a Gaussian on each side that has been raised such that its initial sample is zero. It is parameterized by sigma, which determines the width of the Gaussian rise-fall, the width of the embedded square pulse, and the ratio of the duration of each rise-fall to sigma, in addition to the duration and amplitude.
All native gates on IBM Quantum have predetermined pulses, and calibrations are used periodically to adjust their parameters so they can continue to operate with high fidelity over time.
2.4 Pulse-Level Circuit Description
To completely define a quantum circuit, all necessary pulses must be specified, together with their timing in relation to the circuits’ beginning point and the qubits to which they will be applied. A sequence of pulses (each defined by envelope, frequency, and phase) and the qubits or couplings they operate on effectively defines a so-called pulse-level circuit description. The superconducting quantum computer control equipment generates and delivers the pulses through RF cables to the cryogenic fridge wherein the qubits and couplings are located.
2.5 Running Circuits on Quantum Computers
Figure 4: Example of process for running quantum circuits on superconducting quantum computers in Qiskit.
In Figure 4, we demonstrate a typical Qiskit example on IBM Quantum for running quantum circuits. The quantum circuits are typically represented as code, as shown by QASM Circuit Specification or Qiskit Circuit Specification in Figure 4. Similar to classical computing, quantum circuits typically consist of complex instructions. The preparation, compilation, and assembly processes used for classical computing programs are analogous to the activities needed to convert quantum circuits into low-level and hardware-specific instructions before they can actually be executed on quantum computers. To be more precise, while there can be infinite ways to describe a quantum circuit with the same goal, ultimately only the native gates that are supported by the quantum computer need to be used.
As a result, typically the input circuit specification is translated into a “Gate-Level Circuit", as shown in Figure 4. Gate-level circuits can be visualized as shown in the figure, where the gate operations are represented by the symbols on the lines and qubits are represented by the lines that go from left to right. Without more information, it is usually assumed that qubits are in the |$\ket{0}$ state at the beginning of the quantum circuit. Qubits then evolve through left-to-right sequential processes and are controlled by quantum or classical operations specified in the circuit plot. In order to measure, collect, and store qubit data in classical memory for upcoming analyses, measurements are often carried out at the conclusion of the quantum circuit.
The gate-level circuits are then transpiled, which is a Qiskit term that refers to the operations and transformations that are similar to preprocessing and compilation in classical programs. Transpiling is a multi-step process that involves breaking down non-native quantum gates into groups of native gates, grouping and removing quantum gates to reduce the number of gates, mapping the logic qubits in the original circuits to the physical qubits on the specified quantum computers, routing the circuit under constrained topologies, potentially optimizing circuits to lower error, and more. Following transpilation, circuits are altered in accordance with the knowledge of particular hardware and provide the same logical outcomes as the original circuits. All of the circuits up to this point are gate-level circuits, which employ a more broad description so that they can be executed in many quantum computers. Figure 5 shows one example quantum circuit, and Figure 6 shows one output circuit after transpilation. All the gates are transformed into native gates, and some operations are added to satisfy the topology of the quantum device.
After transpilation, a lower-level procedure occurs, which is known as the schedule in Qiskit. Microwave pulses, which are the final physical processes needed to regulate and control qubits, are further mapped via scheduling to quantum circuits. Due to scheduling, gate-level circuits are converted into pulse-level circuits. The characteristics that define each microwave pulse—such as amplitude, frequency, and others—were previously covered in Section 2.3. Scheduling generates microwave pulse sequences based on calibrated data for each basis gate on each qubit or qubit pair and quantum device. The data includes wave envelopes, frequencies, amplitudes, durations, and other characteristics of microwave pulses. All the information that quantum computers require to run the circuit is contained in the final data. This data will be used to alter the qubits of quantum computers once the quantum circuit has begun, and the qubits themselves are controlled by the equipment.
Using the procedures described above, a set of instructions that may be utilized to carry out the required quantum circuits is created from the original quantum circuits. IBM Quantum offers Qiskit as a tool for users to construct circuits, carry out these actions, and submit quantum circuits to the cloud. The cloud will then carry out the users’ circuits and execute them before returning the results to users.
Figure 5: A quantum adder circuit with width=4 (4 qubits) followed by measurement.
Figure 6: Adder circuit with width=4 transpiled with optimization level 3.
3 Attack Scenario and Threat Model
Figure 7: Attack scenarios based on the attacker’s measurement capabilities.
The operation of modern cloud-based quantum computers allows remote users to submit jobs to the cloud provider. These jobs are dispatched to specific quantum computers, also known as backends. While classical computer components, such as the job management server, are considered protected from side-channels, the controller electronics of quantum computers have not been thoroughly analyzed for potential side-channels. The focus of this work is on potential power trace attacks that could extract information about users’ quantum circuits from the controllers.
3.1 Assumptions of Attacker Measurement
We assume the attacker can sample power traces from shots of a circuit, or they can measure a number of shots and it is easy to divide this into individual shots, since all shots perform the same operations. Recall that each quantum program, i.e. quantum circuit, is executed multiple times, and each execution is called a shot.
We assume the attacker knows when the victim circuits will be executed so the attacker can capture the side-channel information. Precise knowledge of the execution time is not needed as long as the attacker can capture the trace of one shot. Since the victim often executes thousands of shots, the attacker has multiple chances to capture at least one trace. Each shot is identical without considering the noise.
Single-Shot Per-Channel Power Side-Channel Measurement.
A stronger attacker is able to collect per-channel power traces (see Figure 7). The attacker knows directly which pulses are applied to which qubit as each channel controls different qubits and different two-qubit pairs. Such attackers can attempt Circuit Reconstruction attack from per-channel power traces by collecting a single per-channel power trace for each channel.
Single-Shot Total Power Side-Channel Measurement.
A weaker attacker could collect a single total power trace over all channels, but not distinguish the power traces of each channel. Such attackers can attempt Circuit Reconstruction attack by collecting a single total power trace (see Figure 7). In particular, there is a trend to have multiple AWGs being part of same physical device. For example, in QICK [${}^{+}$22] framework, FPGAs are used for waveform generation and one FPGA can generate many control pulses. We believe that going forward total power side-channel attacks may be most realistic as attacker may not easily get power traces for individual channel generated by the FPGA, but can easily measure total power consumption of the FPGA, and thus get total power trace of all the channels. Nevertheless, we explore both per-channel and total power side-channel attacks to understand their potential threats.
3.2 Assumptions of Attacker’s Knowledge
We note that in this work the attacker is assumed to know at all times the information about the target quantum computer (number of qubits it contains, the topology and connections of the qubits) and its basis pulse library. This assumption is reasonable if users have the right to fine-grained control of transpilation and scheduling, because this information is needed in both processes. If this information is not provided, users may easily reverse-engineer it, such as by iteratively increasing the number of qubits to check how many qubits are supported, inserting a two-qubit gate in each qubit pair to check qubit connections, and performing experiments such as frequency sweep and Rabi experiment to acquire the information about the basis pulse library [XES23b].
We assume custom gates are not used by users, and all victim circuits are composed only of the basic gates supported by the quantum computer, typically including ID, RZ, SX, X, and CX for IBM Quantum devices. Among the basic gates, we assume the RZ gates are virtual, as is common today. For an attacker who has only access to collect total power traces, we assume he or she knows the in-channel and cross-channel functions that define how the per-channel and total power traces correspond to the pulse information [XES23b].
3.3 Attacker’s Objective
The attacker aims to uncover quantum circuit details from captured power traces. With access to the basis pulse library, which specifies the pulses for all native gates on a specific quantum device, and the measured power traces of the user’s circuit, the attacker’s goal is to reconstruct the user’s circuit. This means retrieving all necessary information about the user’s circuit to reproduce it. The attacker seeks to learn the transpiled circuit, which is functionally equivalent to the user’s input circuit, even if it may differ in terms of quantum gates used.
3.4 Impact of Attacks
Intellectual property, such as quantum algorithm design, is what many users seek to protect. For instance, proprietary quantum machine learning algorithms are being developed by startups who do not own quantum computers; they are worried about the leakage of their proprietary information. Furthermore, different from classical computing, data in quantum computing is encoded as parts of circuits, such as oracles or ansatzes. Besides, input data such as initial states can also be provided eternally to the execution circuits, but it requires quantum memories and quantum networking, which is not available today. As a result, for example, the circuits used sensitive fields, such as medical-related algorithms, may encode private information, and it needs to be protected.
3.5 The Realism of the Threat Model
Our work focuses on physical side-channel attacks, such as been widely studied in classical computers. As in classical setting, we assume physical access, which is a standard assumption in any physical side-channel attack. The practicality of these attacks is on the same level as for classical computer power side-channel attacks where attackers can probe the power supply network or power supply of the target (signal generator in our case). Note that attackers can purchase signal generators from science equipment vendors to study their power consumption profile and fine-tune attacks ahead of time.
3.6 Difference from Classical Setting of Power Side-Channel Attacks
The major difference of our research is that in classical computers there are no analog control pulses; in classical settings the instructions to the processing unit are digital data read from digital instruction memory, in quantum computers these are analog pulses sent by the signal generators. Our work and threat model assumes any classical and digital information is already protected, and there is a large body of research on the protection of classical computers from power side-channels. Meanwhile, we focus on analog control pulses and signal generators which are not well understood from a security perspective so far.
4 Formalization of Circuit Reconstruction
4.1 Quantum Device
For a superconducting quantum device, sometimes called a quantum processor, the most important features of the topology are the number of qubits and how they are placed and connected with each other. In addition, each quantum device also has its own native gates. In this paper, for a quantum device $D$, we used $n$ to represent its number of qubits, and $m$ to represent its number of qubit connections. The set of basis gates of $D$ is denoted by $BG$. On most of the current quantum devices on IBM Quantum, the basis gates are:
$BG=\{{\tt I,RZ,X,SX,CX}\}.$ (1)
4.2 Channel
Channels refer to which part of the hardware the pulses will be sent to control the qubits. Pulses are applied on one channel for single-qubit gates and several channels for multiple-qubit gates, as described in Section 2.3 on quantum computer controls. The four main categories of channels are drive channels, which send signals to qubits to perform gate operations, control channels, which supplement the drive channel’s control over the qubit, measure channels, which send measurement stimulus pulses for readout, and acquire channels, which are used to gather data. Without considering the measurement operations, quantum circuits only trigger drive and control channels. Drive channels typically correspond to qubits, whereas control channels typically correspond to the connections between qubits and are used for two-qubit gates. The architecture of the quantum computer determines how many channels of each type there are.
For a quantum device $D$ we can define a set $C$ to represent the set of channels on $D$. To be more specific, if only considering drive and control channels, $C$ can be represented as:
$C=\{drive_{0},drive_{1},\dots,drive_{n-1},control_{0},control_{1},\cdots,control_{m-1}\}$ (2)
where $drive$ refers to the drive channel and $control$ refers to the control channel of $D$, and $n$ and $m$ is the number of qubits and connections of the device $D$.
4.3 Basis Pulse
Every quantum circuit must be translated into a quantum circuit that only includes the target quantum device’s basis gates. The group of pulses that follow the scheduling of a basis gate are referred to as its basis pulses. Because the quantum gate is an abstract notion, pulse parameters for the same type of gate on various channels vary because pulse parameters are highly reliant on qubit physical features. For instance, the pulse parameters of the X gate on qubit 00 are often different from those of the X gate on qubits other than 00. Basis gates and their associated pulse waveform are predetermined, thus they typically do not change over different qubits. Thus, to define the basis pulse, the gate type as well as the channels need to be specified. We refer this information to labels, and define the set of labels for all the basis gates and possible channels to be:
$L=\{(gate,C^{\prime})|gate\in BG,C^{\prime}\subset C\}$ (3)
where $BG$ is the set of basis gates and $C$ is the set of channels. $L$ represents all basis gates with their channel information, which can uniquely specify a basis pulse. One basis pulse on the channel $c$ can then be defined as:
$p_{l,c}(x)\begin{cases}\text{Not always 0}&\text{if }x\in[0,d_{l}]\\ =0&\text{if }x\not\in[0,d_{l}]\end{cases}$ (4)
where $c\in l[C^{\prime}],l\in L$ is the label for the basis pulse and $d_{l}$ refers to the duration of the pulse, in discrete time steps. The values of $p_{l,c}(x)$ represent the amplitude of the basis pulse with label $l$ at the channel $c$ on time step $x$. All the time steps are in the unit of the system’s time resolution, which is denoted as dt in Qiskit for IBM Quantum, so the variable $x\in\mathbb{N}$.
Since I, RZ, X, and SX are all single-qubit gates, their basis pulses are made up of only one channel. Whereas, CNOT gate is a two-qubit gate, so it consists of several channels. For most of the quantum devices on IBM Quantum, the duration of single-qubit gates is chosen to be $160$ dt, while the duration of two-qubit gates over different channels is typically different and much longer than the single-qubit gates. For example, $d_{c_{i},{\tt X}}=160$, and $d_{c_{j},{\tt CX}}>1000$ and is often different with different $c_{j}$.
Because one basis pulse may include pulses on several channels, such as CX gate, for each basis gate, its pulses form a set:
$p_{l}(x)=\{p_{l,c}(x)|c\in l[C^{\prime}]\}.$ (5)
4.4 Basis Pulse Library
For all of their quantum devices, IBM Quantum provides the information about basis pulses. We call the collection of basis pulses the basis pulse library. The so-called custom pulse gates, which let users produce their own arbitrary pulses, are another feature supported by IBM Quantum, but are left as future work. Consequently, we assume that there are no custom pulse gates present in the victim circuits. In the end, the basis pulse library can be defined as a set $P_{L}$ which contains all basis pulses.
$P_{L}=\{p_{l}(x)|l\in L\}.$ (6)
4.5 Pulse-Level Circuit
In Section 2, we mentioned a series of instructions describing how to control the qubits with pulses, making reference to the pulse-level circuit. The circuit’s pulse specifications, as well as the start time steps for the instructions, are all contained in the instruction list. One pulse circuit can be formalized as:
$A_{P_{L}}=\left\{a_{l,t}\cdot p_{l}(x-t)|l\in L,a_{l,t}\in\{0,1\},p_{l}(x)\in P_{L}\right\}$ (7)
where its item $a_{l,t}=1$ means that there is a basis pulse of label $l\in L$ being applied which starts from the time step $t$, while $a_{l,t}=0$ means the opposite. As mentioned above, the power traces are discretized in the unit dt, and all the time steps are integers, so $t\in\mathbb{N}$. According to Equation (7), $A_{P_{L}}$ defines all the pulses and where and when they are applied, and thus defines a pulse-level circuit.
4.6 Power Trace
The pulses are generated by classical equipment and thus consume energy. The function of the power value with time is what we refer to as the power trace. The term per-channel power trace refers to the power trace on a single channel, whereas total power trace refers to the function of the summation of power over all channels in a time period. Assume that the ability to monitor power consumption on some or all of the channels exists, and that the measured power trace will be made up of and reliant upon a variety of channels. As it reduces multidimensional data to a single dimension, we refer to the function that creates the total power trace from separate power traces as the summation function or reduction function since it reduces multidimensional data to one-dimensional data.
To formalize the power traces, the per-channel power trace and total power trace functions are needed. The per-channel function $Power_{c}[p_{l}(x)]$, where $c\subseteq C$, specifies how the per-channel power traces are computed. The total power trace function $Total[f_{c}(x)]$, where $c\subseteq C$, specifies how the total power traces are summed up from all per-channel power traces. In the experiment, we assume that the per-channel power traces are the square of the norm of the amplitude:
$Power_{c}[A_{P_{L}}](x)=\sum_{A_{P_{L}}}\text{Re}^{2}[a_{l,t}\cdot p_{l,c}(x-t)]+\text{Im}^{2}[a_{l,t}\cdot p_{l,c}(x-t)]$ (8)
and the total power traces are directly the summation of per-channel power traces:
$Total[A_{P_{L}}](x)=\sum_{c\in C}Power_{c}[A_{P_{L}}](x)$ (9)
4.7 Domain-Specific Constraint
In the circuit, we exploit the following constraint (channel constraint): there can only be at most one pulse on each channel at a given time step, which means:
$\left\{c_{1}=c_{2}\land[t_{1},t_{1}+1,...,t_{1}+d_{gate_{1},c_{1}}]\cap[t_{2},t_{2}+1,...,t_{2}+d_{gate_{2},c_{2}}]=\emptyset\right\}\lor c_{1}\neq c_{2}$ (10)
$\implies\forall t_{1}\ \text{and}\ t_{2}\in[t_{1},t_{1}+1,...,t_{1}+d_{gate_{1},c_{1}}],\ a_{c,gate_{1},t_{1}}+a_{c,gate_{2},t_{2}}\in\{0,1\}$
i.e., if the first pulse with duration $d_{gate_{1},c_{1}}$ has the component on channel $c_{1}$ starting from time step $t_{1}$, and the second pulse with duration $d_{gate_{2},c_{2}}$ has the component on channel $c_{2}$ starting from time step $t_{2}$, then these two components cannot be mixed with each other.
4.8 Attacker’s Goal
For the per-channel single trace attack, the attacker measures $v_{c}(x),\forall c\in C$, the per-channel power traces of the victim circuit. For the total power single trace attack, the attacker measures $v(x)$, the total power traces of the victim circuit. The goal of the attacker is to reconstruct the victim circuit, i.e., find a circuit $A_{P_{L}}$ that is corresponding to the victim circuit. To determine which is better to choose among many circuits, we choose the circuit that minimizes the distance between the power traces with the measured power traces. In addition, the domain-specific constraints discussed in Section 4.7 need to be observed.
For the per-channel single trace attack, the goal is:
$A_{P_{L}}=\operatorname*{arg\,min}_{A^{\prime}_{P_{L}}}\operatorname*{\sum}_{c\in C}{\left(d\left\{Power_{c}[A^{\prime}_{P_{L}}](x),v_{c}(x)\right\}\right)}$ (11)
where $\operatorname*{\sum}_{c\in C}(d_{c})$ is the function to sum up the distances over all channels to get a total distance. This goal is to find the circuit $A_{P_{L}}$ from the set of all circuits $\{A^{\prime}_{P_{L}}\}$ that minimizes the distance between the total power traces of this circuit and the measured per-channel power traces.
For the total power single trace attack, the goal is:
$A_{P_{L}}=\operatorname*{arg\,min}_{A^{\prime}_{P_{L}}}d\left\{Total[A^{\prime}_{P_{L}}](x),v(x)\right\}$ (12)
i.e., finding the circuit $A_{P_{L}}$ from the set of all circuits $\{A^{\prime}_{P_{L}}\}$ that minimizes the distance between the total power traces of this circuit and the measured total power traces.
5 Power Side-Channel Attacks
In this section, we present two methods that we have developed for stealing quantum program secrets. The first method is based on per-channel single trace information, where the attacker uses per-channel measurements to perform a brute-force attack with the goal of reconstructing the quantum program. The second attack is more challenging, as it restricts the attacker to using only a single total power trace to reconstruct the quantum program. Brute-force methods are not scalable in this case, as the sample pulses at each time step are mixed up, as formulated in Equation 12. Therefore, we employ Mixed-Integer Linear Programming optimization to find the set of best pulse-level instructions that decompose the quantum program and their corresponding starting time steps.
In Figure 5, the left part shows a victim circuit (which is a randomly generated for demonstration purposes) that is transpiled on 5-qubit IBM Lima machine (Figure 10(a)). The goal is to recover the circuit from its power trace(s) by finding the set of most suitable pulse-level instructions that make up the circuit with minimum error. The table on the right in Figure 5 shows the starting time step ($\mathrm{dt_{start}}$) of each pulse-level instruction. If the attacker obtains this information, they can reconstruct the circuit, as the order of the instructions is enough to compile the same circuit again. The third column shows a complete quantum program in Python that is used to generate the circuit on the left. While the per-channel attacker aims to recover this table from the measured waveform from each drive channel, the total power attacker employs only a single mixed, superimposed amplitude samples due the the fact that the attacker does not have access to the individual drive channels.
Figure 8: A randomly generated circuit on the left is transpiled on a 5-qubit IBM Lima machine (Figure 10(a)). The table on the right shows the starting index of each pulse-level instruction. We aim to recover this table from the measured waveform from each drive channel for the per-channel attacker, or from the total power trace for the total-power side-channel attacker.
Refer to caption{NiceTabular}|ll|l| $\bf dt_{start}$ Instruction circ = QuantumCircuit(2)
0 SX: $\mathrm{d}_{1}$ circ.
sx(1)
160 SX: $\mathrm{d}_{0}$ circ.
sx(0)
160 SX: $\mathrm{d}_{1}$ circ.
sx(1)
320 CX: $\mathrm{d}_{0}$-$\mathrm{d}_{1}$ circ.
cnot(0, 1)
1696 CX: $\mathrm{d}_{1}$-$\mathrm{d}_{0}$ circ.
cnot(1, 0)
3232 CX: $\mathrm{d}_{0}$-$\mathrm{d}_{1}$ circ.
cnot(0, 1)
5.1 Single-shot Per-Channel Power Side-Channel Attack
The process of reconstructing a quantum circuit in the per-channel single trace attack involves measuring the output of each qubit channel and obtaining the waveform of the pulse sequence from each channel. These waveforms are then compared with a set of pulse sequences that include instructions SX, X, and CX, which are obtained from the pulse library of the quantum computer and are unique to each qubit channel. These pulse sequences can be thought of as profiles in classical side-channel analysis, and the most likely instruction profile needs to be identified. To do that, we compare the measured waveform with each candidate profile and calculate the distance between them. However, measurement errors due to noise, or miscalibration of the measuring device can affect the accuracy of the results. Therefore, it is important to have a reliable means of quantifying the similarity between the measured and candidate waveforms.
In Equation 11, the objective function is defined as the minimum distance between the measured waveform and the candidate pulse sequences. The distance function $d$ can be naturally defined as the Euclidean distance between two power traces:
$d_{2}:\left\{v(x),Total[A_{P_{L}}](x)\right\}\mapsto\|v(x),Total[A_{P_{L}}](x)\|_{2}=\sqrt{\sum_{i=1}^{n}\left(v_{i}(x),Total[A_{P_{L}}](x)\right)^{2}}$ (13)
where $v_{i}(x)$ and $Total_{A_{P}}(x)_{i}$ are the $i$-th elements of $v(x)$ and $Total_{A_{P}}(x)$ respectively. However, we evaluated various distances and metrics on randomly generated circuits including Euclidean distance and found that the Jensen-Shannon distance is the most suitable one for quantifying the candidate instructions against the measurements. This distance metric is commonly used to measure the dissimilarity between probability distributions and is well-suited for the task of comparing the probability distributions of pulse sequences. It has been applied to genome comparison[SJWK09, IHS10] in protein surface comparison [OR03], in machine learning [${}^{+}$20] and particularly in the analysis of the similarity between two quantum states [OBL22]. To compute the Jensen-Shannon distance, we first convert the trace and pulse data of a candidate instruction to two probability vectors $P$ and $Q$ by normalizing the amplitudes to turn them into discrete probability distributions. We then calculate the distance between the two probability vectors by computing the Jensen-Shannon distance (metric) between two probability arrays. This is the square root of the Jensen-Shannon divergence ($\sqrt{{\rm JSD}}$[ES03].
Jensen-Shannon divergence (JSD) is a symmetric, smooth, and bounded measure of dissimilarity between probability distributions that is a well-behaved version of the Kullback-Leibler divergence (KLD) $D_{KL}(P\parallel Q)$. It is widely used in information theory and statistics to measure the distinguishablity between probability distributions. Let $M$ be $\frac{1}{2}(P+Q)$, then the Jensen-Shannon divergence is defined as:
$D_{JS}(P\parallel Q)=\frac{1}{2}D_{KL}(P\parallel M)+\frac{1}{2}D_{KL}(Q\parallel M),$ (14)
For discrete probability distributions $P$ and $Q$ defined on the same sample space, $\mathcal{X}$, the relative entropy from $Q$ to $P$ is defined [Mac03] to be:
$D_{KL}(P\parallel Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right).$ (15)
Over two probability vectors $x$ and $y$, we compute relative entropy, $D_{KL}(P\parallel Q)$, as an elementwise operation as follows:
$D_{KL}(x,y)=\begin{cases}x\log(x/y)&x>0,y>0\\ 0&x=0,y\geq 0\\ \infty&\text{otherwise }\end{cases}$ (16)
Table 1: Different distance measures for each candidate pulse-level instruction against the measured waveform whose error rate is 0.1% while running the circuit in Figure 5 on Drive Channel(1) on IBM Lima machine (Figure 10(a)). Total duration of the circuit is 4608 dt. RMSE: Root Mean Squared Error, $d_{2}$: Euclidean distance, $\sqrt{\rm JSD}$: Jensen-Shannon Divergence distance. For starting indices 160, 320, and 4448 dt, all instructions that can be fit in the time window are considered. While at 160 dt and 320 dt, SX, X and all CX instructions are considered, at 448 dt, only SX and X are considered. $\star$ shows the chosen instruction at a $\text{dt}_{\text{start}}$ whereas $\circ$ indicates a candidate that is considered but not selected. At $\text{dt}=4448$, SX instruction is not selected due to high $\sqrt{\rm JSD}$.
|clclll| $\bf dt_{start}$ $\bf{qubit_{1}}$$\checkmark$ RMSE ${d_{2}}$ $\mathbf{\sqrt{\rm\bf JSD}}$
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
160
CX: $\mathrm{d}_{0}$-$\mathrm{d}_{1}$ 0.00001647 0.0128035 0.402693706
160
CX: $\mathrm{d}_{1}$-$\mathrm{d}_{0}$ 0.00001653 0.0147163 0.414806197
160
CX: $\mathrm{d}_{1}$-$\mathrm{d}_{2}$ 0.00001648 0.0157793 0.420431628
160
CX: $\mathrm{d}_{1}$-$\mathrm{d}_{3}$ 0.00002547 0.0190306 0.433682782
160
CX: $\mathrm{d}_{2}$-$\mathrm{d}_{1}$ 0.00001827 0.0146580 0.428686046
160
CX: $\mathrm{d}_{3}$-$\mathrm{d}_{1}$ 0.00002844 0.0172899 0.391024599
160
SX: $\mathrm{d}_{1}$ normal-$\star$ 0.00000004 0.0000906 0.000000003
160
: $\mathrm{d}_{1}$ 0.00087999 0.0273263 0.000014711
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
320
[…]
320
[…]
320
CX: $\mathrm{d}_{1}$-$\mathrm{d}_{2}$ 0.00000404 0.0078170 0.166908652
320
CX: $\mathrm{d}_{1}$-$\mathrm{d}_{3}$ 0.00002301 0.0180839 0.464903649
320
CX: $\mathrm{d}_{2}$-$\mathrm{d}_{1}$ 0.00000105 0.0035068 0.084267418
320
CX: $\mathrm{d}_{3}$-$\mathrm{d}_{1}$ 0.00002377 0.0158058 0.396308179
320
SX: $\mathrm{d}_{1}$circ 0.00000008 0.0001265 0.000000003
320
: $\mathrm{d}_{1}$ 0.00086605 0.0271090 0.000014711
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
4448
SX: $\mathrm{d}_{2}$ 0.00025480 0.00726239 0.450676207
4448
: $\mathrm{d}_{1}$ 0.00133674 0.03367971 0.450681387
In per-channel single trace attacks, we use the Kullback-Leibler divergence (KLD) and the Jensen-Shannon divergence (JSD) and its distance ($\sqrt{\text{JSD}}$) to distinguish two probability distributions. Therefore, here we provide a brief introduction to divergences, distances and metrics. For more details, we refer the reader to [DDDD09, ${}^{+}$14]. A metric $d$ on a set $\chi$ is a function $d:\chi\times\chi\rightarrow\mathbb{R}_{\geq 0}$ such that for any $x,y,z\in\chi$ the following properties are satisfied:
$\chi$ represents the set of probability distributions and $x$ or $y$ represent an entire probability distribution such as $P=\left\{p_{1},p_{2},\ldots,p_{n}\right\}$ where $p_{i}\geq 0$ and $\forall i,\sum_{i=1}^{n}p_{i}=1$. Often, if a distance measure $d$ only satisfies the property non-negativity, is called a divergence. If, in addition, $d$ satisfies the properties identity of indiscernibles, and symmetry, it is called a distance. Thus, KLD is a divergence, JSD is a distance, and $\sqrt{{\rm JSD}}$ is a metric [OBL22].
Figure 9: Example pulse schedules of SX: $\mathrm{d}_{1}$ ($\star$) and CX: $\mathrm{d}_{0}$-$\mathrm{d}_{1}$ ($\circ$), whose durations are 160dt and 1376dt respectively. As seen from the figure, the first waveform on the drive channel $\mathrm{d}_{1}$ of CX is exactly the same pulse as in SX.
Figure 5 is a randomly generated circuit using SX, X, and CX gates. It is transpiled without optimization over IBM Lima machine in order to keep the layout of the circuit intact for the sake of the presentation. After transpilation, $q_{0}$ maps to Drive Channel 0 and $q_{1}$ maps to Drive Channel 1. The measurement is performed on $q_{0}$ and $q_{1}$.
[…]
We evaluated this metric over the real quantum circuits shown in Table 6 and recovered the correct instruction in all cases under different additive Gaussian noise levels where $\mathcal{N}(x,\sigma)$, that is $x$ is the sample amplitude obtained from power trace and $\sigma$ is the standard deviation varies from 00 (no noise) to $0.1$.